Please help me solve:
An angle measures 14° less than three times its complement.
Find the measure of the angle and it's complement.

Answer by nerdybill(7384)

Two angles are "complementary" if they total to 90 deg.
.
An angle measures 14° less than three times its complement.
Find the measure of the angle and it's complement.
.
Let x = angle
90-x = complement of x
.
x = 3(90-x)-14
x = 270 - 3x -14
x = 256 - 3x
4x = 256
x = 64
.
90-64 = 26
.
The two angles are 64 and 26